OpenStudy (anonymous):

helpo with bernoulli differential equation... y'+xy=xy^2

OpenStudy (divanshu):

what help do you want ? tell full question

OpenStudy (anonymous):

I need to solve using the bernoulli method of u substitution. So far I have u'=-y^(-2)y' and u=y^(-1), substitued as u'-xu=-x

OpenStudy (anonymous):

y' = xy^2 - xy dy / dx = xy^2 - xy dy = (xy^2 - xy) dx dy = (x)(y^2-y) dx dy / (y^2 - y) = x dx

OpenStudy (anonymous):

Someone else continue this. :D

OpenStudy (anonymous):

I realize you can solve it that way, but my prof wants us to use the bernoulli method of u substitution...

OpenStudy (anonymous):

Sorry :'( Don't know